1. When a solution of lead(II) nitrate is mixed with a solution of potassium chromate, a yellow precipitate forms according to the equation.


𝑃𝑏(𝑁𝑂3 )2 (π‘Žπ‘ž) + 𝐾2πΆπ‘Ÿπ‘‚4 (π‘Žπ‘ž) β†’ π‘ƒπ‘πΆπ‘Ÿπ‘‚4 (𝑠) + 2𝐾𝑁𝑂3


(a)What volume of 0.105 M lead (II) nitrate is required to react with 100 mL of 0.120 M potassium chromate

(b)What mass of π‘ƒπ‘πΆπ‘Ÿπ‘‚4 solid forms?


2. Suppose a titration is run in which 25.05 mL of NaOH solution of unknown concentration reacts with 25.00 mL of 0.1000 M 𝐻2𝑆𝑂4 solution. The chemical equation that summarizes the reaction is as follows:


𝐻2𝑆𝑂4 (π‘Žπ‘ž) + 2π‘π‘Žπ‘‚π»(π‘Žπ‘ž) β†’ 2𝐻2𝑂(𝑙) + π‘π‘Ž2𝑆𝑂4(π‘Žπ‘ž)


What is the molarity of the solution?


3. If a solution of 𝐴𝑔𝑁𝑂3 is added to an HCl solution, insoluble AgCl will precipitate:


𝐴𝑔𝑁𝑂3 (π‘Žπ‘ž) + 𝐻𝐢𝑙(π‘Žπ‘ž) β†’ 𝐴𝑔𝐢𝑙(𝑠) + 𝐻𝑁𝑂3(π‘Žπ‘ž)


(a)What volume of 1.20 M 𝐴𝑔𝑁𝑂3 solution must be added to 250.0 mL of a 0.100 M HCl solution to precipitate all the chloride ion?

(b)What mass of AgCl should precipitate?


1
Expert's answer
2022-04-13T04:42:03-0400

1.) 𝑃𝑏(𝑁𝑂3Β )2Β (π‘Žπ‘ž) + 𝐾2πΆπ‘Ÿπ‘‚4Β (π‘Žπ‘ž) β†’ π‘ƒπ‘πΆπ‘Ÿπ‘‚4Β (𝑠) + 2𝐾𝑁𝑂3;

C(𝑃𝑏(𝑁𝑂3Β )2) = 0.105 M;

C(𝐾2πΆπ‘Ÿπ‘‚4) = 0.120 M;

V(𝐾2πΆπ‘Ÿπ‘‚4) = 100 mL = 0.100 L;

a) C1 * V1 = C2 * V2

V(𝑃𝑏(𝑁𝑂3Β )2) = C(𝐾2πΆπ‘Ÿπ‘‚4) * V(𝐾2πΆπ‘Ÿπ‘‚4)/C(𝑃𝑏(𝑁𝑂3Β )2) = 0.120 * 0.100/0.105 = 0.114 L = 114 mL;

b) n(𝐾2πΆπ‘Ÿπ‘‚4) = C(𝐾2πΆπ‘Ÿπ‘‚4) * V(𝐾2πΆπ‘Ÿπ‘‚4) = 0.120 * 0.1 = 0.012 mol;

n(π‘ƒπ‘πΆπ‘Ÿπ‘‚4) = n(𝐾2πΆπ‘Ÿπ‘‚4) = n(𝑃𝑏(𝑁𝑂3Β )2) = 0.012 mol;

M(π‘ƒπ‘πΆπ‘Ÿπ‘‚4) = 323 g/mol;

m(π‘ƒπ‘πΆπ‘Ÿπ‘‚4) = n(π‘ƒπ‘πΆπ‘Ÿπ‘‚4) * M(π‘ƒπ‘πΆπ‘Ÿπ‘‚4) = 0.012 * 323 = 3.88 g.



2.) 𝐻2𝑆𝑂4Β (π‘Žπ‘ž) + 2π‘π‘Žπ‘‚π»(π‘Žπ‘ž) β†’ 2𝐻2𝑂(𝑙) + π‘π‘Ž2𝑆𝑂4(π‘Žπ‘ž)

V(H2SO4) = 25.00 mL = 0.02500 L;

C(H2SO4) = 0.1000 M;

V(NaOH) = 25.05 mL = 0.02505 L;

C(H2SO4) * V(H2SO4) = C(NaOH) * V(NaOH)/2;

C(NaOH) = C(H2SO4) * V(H2SO4) * 2/V(NaOH) = 0.1000 * 25.00 * 2/25.05 = 0.1996 M.


3.) 𝐴𝑔𝑁𝑂3Β (π‘Žπ‘ž) + 𝐻𝐢𝑙(π‘Žπ‘ž) β†’ 𝐴𝑔𝐢𝑙(𝑠) + 𝐻𝑁𝑂3(π‘Žπ‘ž)

V(HCl) = 250.0 mL = 0.2500 L;

C(HCl) = 0.100 M;

C(𝐴𝑔𝑁𝑂3) = 1.200 M

a) C(𝐴𝑔𝑁𝑂3) * V(𝐴𝑔𝑁𝑂3) = C(HCl) * V(HCl);

V(𝐴𝑔𝑁𝑂3) = C(HCl) * V(HCl)/C(𝐴𝑔𝑁𝑂3) = 0.100 * 0.2500/1.2 = 0.0208 = 20.8 mL.

b) n(HCl) = C(HCl) * V(HCl) = 0.100 * 0.2500 = 0.025 mol;

n(AgCl) = n(HCl) = n(𝐴𝑔𝑁𝑂3) = 0.025 mol;

M(AgCl) = 143 g/mole

m(AgCl) = n(AgCl) * M(AgCl) = 0.025 * 143 = 3.6 g.